# Generating Multi-Table Time Series EHR from Latent Space with Minimal Preprocessing

### v1

@@ Abstract
Electronic Health Records (EHR) are time-series relational databases that record patient interactions and medical events over time, serving as a critical resource for healthcare research and applications. However, privacy concerns and regulatory restrictions limit the sharing and utilization of such sensitive data, necessitating the generation of synthetic EHR datasets. Unlike previous EHR synthesis methods—which typically generate medical records consisting of expert-chosen features (e.g., a few vital signs, structured codes only)—we introduce RawMed, the first framework to synthesize multi-table, time-series EHR data that closely resembles raw EHRs. Using text-based representation and compression techniques, RawMed captures complex structures and temporal dynamics with minimal preprocessing. We also propose a new evaluation framework for multi-table time-series synthetic EHRs, assessing distributional similarity, inter-table relationships, temporal dynamics, and privacy. Validated on two open-source EHR datasets, RawMed outperforms baseline models in fidelity and utility. The code is available at https://github.com/eunbyeol-cho/RawMed.
@@ 1 Introduction
The digitalization of medical data has accelerated the adoption of Electronic Health Records (EHR), one of the most significant innovations in modern healthcare. EHRs systematically store various medical events, such as prescriptions and test results, along with corresponding timestamps in a multi-table relational database, encompassing categorical, numerical, and text data. Due to these properties, EHRs serve as a crucial resource in various medical AI studies, including clinical prediction, information retrieval, and question answering Goldstein2017OpportunitiesAC ; Shickel18DeepEHR .
@@
However, since EHRs contain sensitive personal information, they are subject to privacy regulations that hinder data sharing and research applications Benitez2010EvaluatingRR . In response, synthetic data has emerged as a promising solution, offering a privacy-preserving alternative that can still support research and clinical applications yan2022multifaceted ; gonzales2023synthetic . Progress in generative models, especially Generative Adversarial Networks (GANs) gan , has opened up new avenues for synthesizing EHR data. Early efforts primarily generated discrete features or heterogeneous features, often overlooking the time-series aspect Choi2017GeneratingMD ; torfi2020corgancorrelationcapturingconvolutionalgenerative . More recent studies incorporate both time-series and heterogeneous features, gradually approaching a closer reflection of real EHR data emrmgan ; ehrsafe ; flexgenehr ; timediff .
@@
Despite this progress, existing EHR synthesis studies still face the following limitations. First, there is a high dependence on feature selection. Most approaches emrmgan ; ehrsafe ; timediff typically select only a subset of tables and columns of the entire EHR database based on domain knowledge, and generate synthetic data exclusively for the features within those selected columns. [^1] In this paper, a column denotes a raw data field in an EHR table, while a feature refers to a processed data representation derived from one or more columns. See Appendix A.1. While this method simplifies the synthesis process, it limits the usefulness of the generated data. For example, if a new research question or analysis requires a feature that was excluded, it is impossible to build predictive models or perform statistical analyses based on that synthetic EHR data. Morever, prior research fullfeature1 ; fullfeature2 ; genhpf suggests that models incorporating a broader range of features tend to achieve higher predictive accuracy, implying that richer synthetic data offers wider utility and more substantial value across various downstream tasks. Second, many existing approaches rely on complex preprocessing steps. Techniques such as numeric binning, one-hot encoding, term normalization, and aggregation are commonly used, but can unintentionally distort the data or obscure meaningful patterns. For instance, aggregating lab values over time may mask sudden anomalies while binning them into discrete ranges may oversimplify subtle trends, reducing the synthetic data’s fidelity for predictive modeling.
@@
emrmgan 98 24 $\times$ $\times$
ehrsafe 90 50 $\times$ $\times$
flexgenehr 5,373 48 $\times$ $\times$
timediff 15 276 $\times$ $\times$
@@
To address these challenges, our framework, RawMed, synthesizes raw EHRs, multi-table time-series data that preserve all columns and original values in their database form, as illustrated in Figure 1. To implement this approach, RawMed adopts a text-based method, treating EHR data as text to retain original values without additional transformations (e.g., binning, one-hot encoding, normalization). This not only reduces the risk of data distortion, but also makes our method generalizable to various EHR data with different database schemas. However, as time-series data becomes longer, its textual representation typically grows even longer due to subword tokenization, leading to a significant increase in computational complexity.
To mitigate this, we employ Residual Quantization rqvae to compress textualized time-series data, enabling autoregressive modeling in a compressed latent space. This reduces sequence length and computational complexity, allowing RawMed to handle multi-table EHR datasets with numerous columns. Consequently, our framework generates raw EHRs with extensive features and time steps surpassing most existing methods (Table 1). Our key contributions are summarized as follows:
@@ 2.1 EHR data generation
Early research on EHR synthesis primarily focused on generating single data types (e.g., diagnostic codes) using generative models such as GANs and VAEs. Subsequent studies explored mixed data types or incorporated time-series continuity esteban2017real ; biswal2020eva ; zhang2021synteg ; wang2022promptehrconditionalelectronichealthcare ; Theodorou_2023 . However, relatively few studies have addressed both the temporal dynamics and heterogeneous nature of EHR data simultaneously, a gap that has only recently started to gain attention. For example, EMR-M-GAN emrmgan pioneered a GAN-based approach for mixed-type time-series EHR data, using a Dual-VAE and Coupled Recurrent Network (CRN) to capture inter-feature correlations and temporal dynamics. EHR-Safe ehrsafe also adopted a GAN-based framework designed to handle real-world EHR complexities comprehensively, including missing data, diverse feature types, and both static and time-varying attributes. FLEXGEN-EHR flexgenehr integrated static and time-series data, focusing on missing values via optimal transport. TIMEDIFF timediff introduced a hybrid diffusion model capable of generating both continuous and discrete time-series data concurrently. To the best of our knowledge, no existing studies do not utilize all columns from the original EHR database tables. Instead, they rely on a subset of features deemed important by the researchers, and heavily preprocess (e.g., numerical binning, one-hot encoding, normalization, aggregation) the original EHR data to train the generative model.
@@ 2.2 Text-based approaches for tabular data generation
With the rapid progress of NLP, studies have emerged that convert tabular data into textual form for generation purposes great ; tabula . This text-based approach offers several advantages: it minimizes complex preprocessing, preserves original values without information loss, and leverages the pre-trained knowledge of language models. Such benefits are particularly evident for datasets like EHRs, which often follow varying database schemas and require substantial preprocessing. Nevertheless, most existing research on tabular data synthesis focuses on single-table settings great ; tabula , with comparatively few studies addressing relational database realtabformer and no studies synthesizing time-series relational database. Moreover, converting tabular data into text increases the length of each row due to subword tokenization, potentially inflating training and inference costs. To mitigate this, RawMed builds on text-based representations but compresses textualized data, enabling efficient autoregressive modeling in a latent space. This approach reduces computational cost while preserving the fidelity of raw, multi-table time-series EHRs across diverse database schemas.
@@ 3 Method
RawMed is a framework for synthesizing multi-table time-series EHRs, capturing complex patient trajectories with minimal preprocessing. Its architecture supports heterogeneous data type and varying event counts using text-based event representation, with two modules: (1) event-level compression for compact event encoding and (2) temporal inter-event modeling for dynamic temporal relationships. This design enables precise synthesis of large-scale raw EHR datasets.
@@ 3.1 Data representation and notation
[…]
@@ 3.2 Compressing event representations
To address this, RawMed compress the event text embedding $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$, where $L$ is the number of tokens and $F$ is the embedding dimension, into a discrete latent representation $\mathbf{z}_{i}^{p}\in\mathbb{R}^{L_{z}\times F_{z}}$, where $L_{z}$ is the latent sequence length and $F_{z}$ is the latent embedding dimension. We adopt neural network-based compression methods, Vector Quantized Variational AutoEncoders (VQ-VAE) vqvae and Residual Quantization (RQ) rqvae , implemented with 1D convolutional neural networks (CNNs).
@@
The encoder Enc transforms $\mathbf{x}_{i}^{p}$ into $\hat{\mathbf{z}}_{i}^{p}=\texttt{Enc}(\mathbf{x}_{i}^{p})\in\mathbb{R}^{L_{z}\times F_{z}}$, consisting of $L_{z}$ latent vectors. Each latent vector, denoted $\hat{\mathbf{z}}\in\mathbb{R}^{F_{z}}$ for simplicity, is quantized to $\mathbf{z}\in\mathbb{R}^{F_{z}}$ by assigning it to the nearest entry in the codebook $C=\{(k,\texttt{lut}(k))\mid k=1,\dots,K\}$, where $\texttt{lut}(k)\in\mathbb{R}^{F_{z}}$ is the embedding for index $k$ [^2] The function lut refers to a look-up table that maps indices to their corresponding embeddings. The quantization is defined as:
@@
where $D$ is the quantization depth (i.e., $D$=1 for VQ). The RQ process initializes the residual vector as $\mathbf{r}_{0}=\hat{\mathbf{z}}$. For each depth $d=1,\dots,D$, it quantizes the residual to obtain an index $k_{d}=\texttt{VQ}(\mathbf{r}_{d-1};C)$ and computes the next residual as $\mathbf{r}_{d}=\mathbf{r}_{d-1}-\texttt{lut}(k_{d})$. Partial sums $\mathbf{z}^{(d)}=\sum_{m=1}^{d}\texttt{lut}(k_{m})$ yield increasingly refined approximations, with $\mathbf{z}=\mathbf{z}^{(D)}$.
@@ 3.4 Data sampling and postprocessing
Synthetic patient trajectories are generated by autoregressively sampling token sequences using the trained TempoTransformer with Top-$k$ sampling. Each synthetic sequence $\tilde{S}^{p}$ interleaves time tokens $\tau_{i}^{p}$ and event tokens $k_{i}^{p}$. Sampling is constrained to ensure structural integrity: the first $L_{t}$ tokens per block, representing time, are sampled from the time token vocabulary $[0,9]$, while the subsequent $L_{z}\cdot D$ event tokens are sampled from the event token vocabulary $[K]$, with invalid token probabilities masked. Next, the event tokens are mapped to latent representations $\tilde{\mathbf{z}}_{i}^{p}$ via codebook $C$, decoded by Dec into text embeddings $\tilde{\mathbf{x}}_{i}^{p}$, and converted to text (e.g., lab item Glucose value 95 uom mg/dL). Time tokens are detokenized into timestamps (e.g., $[7,2]\to 720~{}\text{minutes}$). These text-based events and timestamps are then converted to relational tables.
@@ 4.1 Setup
In this study, we used two publicly available EHR datasets, MIMIC-IV mimiciv and eICU eicu . Adhering to the GenHPF genhpf framework, which utilizes all EHR features for diverse prediction tasks, we focused on patients aged 18 years or older who had been admitted to the ICU for more than 12 hours. Based on this framework, we integrated three time-series tables (laboratory tests, prescriptions, and input events) from both datasets. To assess the generalizability of our approach across diverse clinical settings, we conducted additional experiments on the MIMIC-IV-ED mimicived emergency department dataset, with results reported in Appendix F.3. After preprocessing, the datasets comprised 62,704 patients (about 7 million rows) from MIMIC-IV and 143,812 (about 7 million rows) from eICU, with summary statistics detailed in Table 7. The dataset was split into a 9:1 ratio, allocating 90% for training and 10% for testing. The generative model was trained exclusively with the training data. We aimed to generate data across all columns to preserve the original dataset as comprehensively as possible; however, we excluded columns offering no analytical value (e.g., subject_id, orderid). Details on exclusions are provided in Appendix B.3.
@@
This study is the first to synthesize raw EHR data, and therefore does not share an identical problem formulation with prior EHR synthesis research, making direct comparisons challenging. Because existing methods rely on predetermined feature selection and complex preprocessing steps, it is infeasible to extend them to the broader generative scope pursued in this work (i.e., all elements in the original database structure). To provide a baseline, we adapt RealTabFormer realtabformer , designed for text-based relational databases rather than time-series relational databases. Built on GPT-2, RealTabFormer is constrained by limited context length, restricting the amount of data it can process. To address this, we concatenate time-series data in chronological order and apply QLoRA dettmers2024qlora to Llama 3.1dubey2024llama , rather than GPT-2 gpt2 , allowing for longer sequences. Moreover, RealTabFormer does not compress data but instead models it directly, effectively serving as a no-compression baseline.
Additionally, we include baselines using SDV sdv (Gaussian Copula-based), RC-TGAN rctgan (GAN-based), and ClavaDDPM clavaddpm (diffusion-based) for multi-table generation. These methods, while not designed for time-series multi-table data, can generate synthetic data by modeling timestamps as standard columns within each table, as is typical in EHR database structures. However, these approaches might struggle to capture the temporal dynamics due to their lack of explicit temporal modeling. Appendix E.1 details the implementation and technical specifications.

### v2

# arXiv:2507.06996v2[cs.LG] 02 Mar 2026

@@ Abstract
Electronic Health Records (EHR) are time-series relational databases that record patient interactions and medical events over time, serving as a critical resource for healthcare research and applications. However, privacy concerns and regulatory restrictions limit the sharing and utilization of such sensitive data, necessitating the generation of synthetic EHR datasets. Unlike previous EHR synthesis methods—which typically generate medical records consisting of expert-chosen features (e.g., a few vital signs, structured codes only)—we introduce RawMed, the first framework to synthesize multi-table, time-series EHR data that closely resembles raw EHRs. Using text-based representation and compression techniques, RawMed captures complex structures and temporal dynamics with minimal lossy preprocessing. We also propose a new evaluation framework for multi-table time-series synthetic EHRs, assessing distributional similarity, inter-table relationships, temporal dynamics, and privacy. Validated on two open-source EHR datasets, RawMed outperforms baseline models in fidelity and utility. The code is available at https://github.com/eunbyeol-cho/RawMed.
@@ 1 Introduction
The digitalization of medical data has accelerated the adoption of Electronic Health Records (EHR), one of the most significant innovations in modern healthcare. EHRs systematically store various medical events, such as prescriptions and test results, along with corresponding timestamps in a multi-table relational database, encompassing categorical, numerical, and text data. Due to these properties, EHRs serve as a crucial resource in various medical AI studies, including clinical prediction, information retrieval, and question answering Goldstein et al. (2017); Shickel et al. (2018).
@@
However, since EHRs contain sensitive personal information, they are subject to privacy regulations that hinder data sharing and research applications Benitez and Malin (2010). In response, synthetic data has emerged as a promising solution, offering a privacy-preserving alternative that can still support research and clinical applications Yan et al. (2022); Gonzales et al. (2023). Progress in generative models, especially Generative Adversarial Networks (GANs) Goodfellow et al. (2014), has opened up new avenues for synthesizing EHR data. Early efforts primarily generated discrete features or heterogeneous features, often overlooking the time-series aspect Choi et al. (2017); Torfi and Fox (2020). More recent studies incorporate both time-series and heterogeneous features, gradually approaching a closer reflection of real EHR data Li et al. (2023); Yoon et al. (2023); He et al. (2024); Tian et al. (2024).
@@
Despite this progress, existing EHR synthesis studies still face the following limitations. First, there is a high dependence on feature selection. Most approaches Li et al. (2023); Yoon et al. (2023); Tian et al. (2024) typically select only a subset of tables and columns of the entire EHR database based on domain knowledge, and generate synthetic data exclusively for the features within those selected columns. [^1] In this paper, a column denotes a raw data field in an EHR table, while a feature refers to a processed data representation derived from one or more columns. See Appendix A.1. While this method simplifies the synthesis process, it limits the usefulness of the generated data. For example, if a new research question or analysis requires a feature that was excluded, it is impossible to build predictive models or perform statistical analyses based on that synthetic EHR data. Morever, prior research Rajkomar et al. (2018); Nallabasannagari et al. (2020); Hur et al. (2023) suggests that models incorporating a broader range of features tend to achieve higher predictive accuracy, implying that richer synthetic data offers wider utility and more substantial value across various downstream tasks. Second, many existing approaches rely on complex, lossy preprocessing steps. Techniques such as numeric binning, term normalization, and aggregation are commonly used, but can unintentionally distort the data or obscure meaningful patterns. For instance, aggregating lab values over time may mask sudden anomalies while binning them into discrete ranges may oversimplify subtle trends, reducing the synthetic data’s fidelity for predictive modeling.
@@
Li et al. (2023) 98 24 $\times$ $\times$
Yoon et al. (2023) 90 50 $\times$ $\times$
He et al. (2024) 5,373 48 $\times$ $\times$
Tian et al. (2024) 15 276 $\times$ $\times$
@@
To address these challenges, our framework, RawMed, synthesizes raw EHRs, multi-table time-series data that preserve all columns and original values in their database form, as illustrated in Figure 1. To implement this approach, RawMed adopts a text-based method, treating EHR data as text to retain original values without additional transformations (e.g., binning, aggregation, or term normalization). This not only reduces the risk of data distortion, but also makes our method generalizable to various EHR data with different database schemas. However, as time-series data becomes longer, its textual representation typically grows even longer due to subword tokenization, leading to a significant increase in computational complexity.
To mitigate this, we employ Residual Quantization Lee et al. (2022) to compress textualized time-series data, enabling autoregressive modeling in a compressed latent space. This reduces sequence length and computational complexity, allowing RawMed to handle multi-table EHR datasets with numerous columns. Consequently, our framework generates raw EHRs with extensive features and time steps surpassing most existing methods (Table 1). Our key contributions are summarized as follows:
@@ 2.1 EHR data generation
Early research on EHR synthesis primarily focused on generating single data types (e.g., diagnostic codes) using generative models such as GANs and VAEs. Subsequent studies explored mixed data types or incorporated time-series continuity Esteban et al. (2017); Biswal et al. (2020); Zhang et al. (2021); Wang and Sun (2022); Theodorou et al. (2023b, a). However, relatively few studies have addressed both the temporal dynamics and heterogeneous nature of EHR data simultaneously, a gap that has only recently started to gain attention. For example, EMR-M-GAN Li et al. (2023) pioneered a GAN-based approach for mixed-type time-series EHR data, using a Dual-VAE and Coupled Recurrent Network (CRN) to capture inter-feature correlations and temporal dynamics. EHR-Safe Yoon et al. (2023) also adopted a GAN-based framework designed to handle real-world EHR complexities comprehensively, including missing data, diverse feature types, and both static and time-varying attributes. FLEXGEN-EHR He et al. (2024) integrated static and time-series data, focusing on missing values via optimal transport. TIMEDIFF Tian et al. (2024) introduced a hybrid diffusion model capable of generating both continuous and discrete time-series data concurrently. To the best of our knowledge, no existing studies utilize all columns from the original EHR database tables. Instead, they rely on a subset of features deemed important by the researchers, and heavily preprocess (e.g., numerical binning, aggregation) the original EHR data to train the generative model.
@@ 2.2 Text-based approaches for tabular data generation
With the rapid progress of NLP, studies have emerged that convert tabular data into textual form for generation purposes Borisov et al. (2022); Zhao et al. (2023). This text-based approach offers several advantages: it minimizes complex preprocessing, preserves original values without information loss, and leverages the pre-trained knowledge of language models. Such benefits are particularly evident for datasets like EHRs, which often follow varying database schemas and require substantial preprocessing. Nevertheless, most existing research on tabular data synthesis focuses on single-table settings Borisov et al. (2022); Zhao et al. (2023), with comparatively few studies addressing relational database Solatorio and Dupriez (2023) and no studies synthesizing time-series relational database. Moreover, converting tabular data into text increases the length of each row due to subword tokenization, potentially inflating training and inference costs. To mitigate this, RawMed builds on text-based representations but compresses textualized data, enabling efficient autoregressive modeling in a latent space. This approach reduces computational cost while preserving the fidelity of raw, multi-table time-series EHRs across diverse database schemas.
@@ 3 Method
RawMed is a framework for synthesizing multi-table time-series EHRs, capturing complex patient trajectories with minimal lossy preprocessing. Its architecture supports heterogeneous data type and varying event counts using text-based event representation, with two modules: (1) event-level compression for compact event encoding and (2) temporal inter-event modeling for dynamic temporal relationships. This design enables precise synthesis of large-scale raw EHR datasets.
@@ 3.1 Data representation and notation
To create a textual representation, we serialize each event by concatenating the table name $\epsilon_{i}^{p}$, followed by column names and their values, excluding null-valued columns. For example, an event from the lab table with $a_{i}^{p}=\{(\text{item},\text{``Glucose''}),\ (\text{value},\ \text{``95''}),\ (\text{uom},\ \text{``mg/dL''})\}$ is serialized as “lab item Glucose value 95 uom mg/dL”. Thus, an event is denoted as $e_{i}^{p}=(t_{i}^{p},x_{i}^{p})$, where $x_{i}^{p}$ is the serialized text string. Then $x_{i}^{p}$ is tokenized, padded or truncated to a fixed length $L$ (typically 128), and embedded to $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$. Given the high dimensionality of these embeddings, compression is essential to model the entire patient sequence $S^{p}$, a challenge also addressed in the context of text-based EHR embeddings Cho et al. (2023).
@@ 3.2 Compressing event representations
To address this, RawMed compress the event text embedding $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$, where $L$ is the number of tokens and $F$ is the embedding dimension, into a discrete latent representation $\mathbf{z}_{i}^{p}\in\mathbb{R}^{L_{z}\times F_{z}}$, where $L_{z}$ is the latent sequence length and $F_{z}$ is the latent embedding dimension. We adopt neural network-based compression methods, Vector Quantized Variational AutoEncoders (VQ-VAE) Van Den Oord et al. (2017) and Residual Quantization (RQ) Lee et al. (2022), implemented with 1D convolutional neural networks (CNNs).
@@
The encoder Enc transforms $\mathbf{x}_{i}^{p}$ into $\hat{\mathbf{z}}_{i}^{p}=\texttt{Enc}(\mathbf{x}_{i}^{p})\in\mathbb{R}^{L_{z}\times F_{z}}$, consisting of $L_{z}$ latent vectors. Each latent vector, denoted $\hat{\mathbf{z}}\in\mathbb{R}^{F_{z}}$ for simplicity, is quantized to $\mathbf{z}\in\mathbb{R}^{F_{z}}$ by assigning it to the nearest entry in the codebook $C=\{(k,\texttt{lut}(k))\mid k=1,\dots,K\}$, where $\texttt{lut}(k)\in\mathbb{R}^{F_{z}}$ is the embedding for index kk. [^2] The function lut refers to a look-up table that maps indices to their corresponding embeddings. The quantization is defined as:
@@
where $D$ is the quantization depth (i.e., DD=1 for VQ). The RQ process initializes the residual vector as $\mathbf{r}_{0}=\hat{\mathbf{z}}$. For each depth $d=1,\dots,D$, it quantizes the residual to obtain an index $k_{d}=\texttt{VQ}(\mathbf{r}_{d-1};C)$ and computes the next residual as $\mathbf{r}_{d}=\mathbf{r}_{d-1}-\texttt{lut}(k_{d})$. Partial sums $\mathbf{z}^{(d)}=\sum_{m=1}^{d}\texttt{lut}(k_{m})$ yield increasingly refined approximations, with $\mathbf{z}=\mathbf{z}^{(D)}$.
@@ 3.4 Data sampling and postprocessing
Synthetic patient trajectories are generated by autoregressively sampling token sequences using the trained TempoTransformer with Top-$k$ sampling. Each synthetic sequence $\tilde{S}^{p}$ interleaves time tokens $\tau_{i}^{p}$ and event tokens $k_{i}^{p}$. Sampling is constrained to ensure structural integrity: the first $L_{t}$ tokens per block, representing time, are sampled from the time token vocabulary $[0,9]$, while the subsequent $L_{z}\cdot D$ event tokens are sampled from the event token vocabulary $[K]$, with invalid token probabilities masked. Next, the event tokens are mapped to latent representations $\tilde{\mathbf{z}}_{i}^{p}$ via codebook $C$, decoded by Dec into text embeddings $\tilde{\mathbf{x}}_{i}^{p}$, and converted to text (e.g., lab item Glucose value 95 uom mg/dL). Time tokens are detokenized into timestamps (e.g., $7,2]\to 720~\text{minutes}$). These text-based events and timestamps are then converted to relational tables.
@@ 4.1 Setup
In this study, we used two publicly available EHR datasets, MIMIC-IV Johnson et al. (2020) and eICU Pollard et al. (2018). Adhering to the GenHPF Hur et al. (2023) framework, which utilizes all EHR features for diverse prediction tasks, we focused on patients aged 18 years or older who had been admitted to the ICU for more than 12 hours. Based on this framework, we integrated three time-series tables (laboratory tests, prescriptions, and input events) from both datasets. To assess the generalizability of our approach across diverse clinical settings, we conducted additional experiments on the MIMIC-IV-ED Johnson et al. (2023) emergency department dataset, with results reported in Appendix F.3. After preprocessing, the datasets comprised 62,704 patients (about 7 million rows) from MIMIC-IV and 143,812 (about 7 million rows) from eICU, with summary statistics detailed in Table 7. The dataset was split into a 9:1 ratio, allocating 90% for training and 10% for testing. The generative model was trained exclusively with the training data. We aimed to generate data across all columns to preserve the original dataset as comprehensively as possible; however, we excluded columns offering no analytical value (e.g., subject_id, orderid). Details on exclusions are provided in Appendix B.3.
@@
This study is the first to synthesize raw EHR data, and therefore does not share an identical problem formulation with prior EHR synthesis research, making direct comparisons challenging. Because existing methods rely on predetermined feature selection and complex preprocessing steps, it is infeasible to extend them to the broader generative scope pursued in this work (i.e., all elements in the original database structure). To provide a baseline, we adapt RealTabFormer Solatorio and Dupriez (2023), designed for text-based relational databases rather than time-series relational databases. Built on GPT-2, RealTabFormer is constrained by limited context length, restricting the amount of data it can process. To address this, we concatenate time-series data in chronological order and apply QLoRA Dettmers et al. (2024) to Llama 3.1Dubey et al. (2024), rather than GPT-2 Radford et al. (2019), allowing for longer sequences. Moreover, RealTabFormer does not compress data but instead models it directly, effectively serving as a no-compression baseline.
Additionally, we include baselines using SDV Patki et al. (2016) (Gaussian Copula-based), RC-TGAN Gueye et al. (2023) (GAN-based), and ClavaDDPM Pang et al. (2024) (diffusion-based) for multi-table generation. These methods, while not designed for time-series multi-table data, can generate synthetic data by modeling timestamps as standard columns within each table, as is typical in EHR database structures. However, these approaches might struggle to capture the temporal dynamics due to their lack of explicit temporal modeling. Appendix E.1 details the implementation and technical specifications.